Shanette paid 20 pesos for materials to make earnings. After working hard to make them, she made a total of 10 earnings. She then sold 7 of the earning for 5 pesos each, and 3 others for 2 pesos each. How much is shanette's profit?

1
Expert's answer
2021-12-20T16:05:01-0500

Total pesos earned;

7×5 =35

3×2=6

35+6=41

profit = 41-20=21pesos